OpenStudy (anonymous):

How many solutions does the following system of equations have? y=5/2x+2 2y=5x+4 One Two Zero Infinitely many

Directrix (directrix):

Take this equation --> y=5/2x+2 and multiply each term by 2. Post what you get.

Directrix (directrix):

@mikeshaw

OpenStudy (anonymous):

y=9x+4

OpenStudy (anonymous):

Is that right? @Directrix

OpenStudy (anonymous):

anyone help me?

OpenStudy (mandre):

Multiply each term by 2 as Directrix said: What is y * 2? What is 5/2x * 2?

OpenStudy (mandre):

How did you get 9??

OpenStudy (anonymous):

2y=5x+4*

OpenStudy (mandre):

There you go. Now you have : 2y=5x+4 2y=5x+4 You have 2 identical equations so you actually only have 1 equation: 2y=5x+4 What can you deduce from that about your question?

OpenStudy (anonymous):

That you can only have one solution?

hartnn (hartnn):

or you can do one thing, subtract the first equation from second what u get ?

OpenStudy (mandre):

What if x = -1 and y = -1/2? What if x = 1 and y = 9/2?

OpenStudy (mandre):

You still think there's only one solution?

OpenStudy (anonymous):

So there is many solutions then

OpenStudy (mandre):

So your answer is?

OpenStudy (anonymous):

infinetly many

OpenStudy (mandre):

There you go. Whenever you have one equation but more than one unkown like 2x + y = 10 then you will always have infinitely many solutions.

OpenStudy (anonymous):

tjhank you :D

OpenStudy (mandre):

Pleasure.